slove the follwing system of equation using subatitution 3x+6y=48 and-5x+6y=32
Answer by checkley77(12844)

3x+6y=48
6y=-3x+48 eq#1
-5x+6y=32
6y=5x+32 eq#2
set these 2 equations equal
-3x+48=5x+32
-3x-5x=32-48
-8x=-16
x=-16/-8
x=2 ans.
3*2+6y=48
6+6y=48
6y=48-6
6y=42
y=52/6
y=7 ans.
Proof:
-5*2+6*7=32
-10+42=32
32=32